OpenStudy (yinshy):

Crestline Bank offers 3% simple interest on deposits. Mid City Bank offers 2% interest compounded annually. Jill deposits $3000 in Crestline Bank and Owen deposits $3000 in Mid City Bank. What is the difference in final balances for Jill and Owen at the end of 5 years? A. $177.82 B. $165.58 C. $150.00 D. $137.76

OpenStudy (greatlife44):

One is compounded yearly hence annually while the other is just simple interest.

OpenStudy (yinshy):

so A?

OpenStudy (mathstudent55):

Have you done the calculations to think the answer is A? If so, show them.

OpenStudy (yinshy):

is it A?

OpenStudy (greatlife44):

For the first one we can use the simple interest formula for the account in crestline bank where p = principle starting value, r is the interest rate and t = time in years. \[i = p*r*t\] \[(3*10^{3})(\frac{ 3 }{ 100 })(5) = $450\] for the other one we need to take into account that the interest is compounded. annualy.

OpenStudy (greatlife44):

Compound interest formula where P is your starting value, r is your interest rate, and n is the number of times your interest rate is compounded. and t is in years. \[A = P(1+\frac{ r }{n })^{nt}\]

OpenStudy (greatlife44):

follow?

OpenStudy (yinshy):

kinda not the second one

OpenStudy (greatlife44):

BTW that formula gives the interest gained so I think it's actually 450+3,000 our initial deposit. 3450 is what we end up getting.

OpenStudy (greatlife44):

P is our principle 3,000 r is our interest rate which is 2/100 0.02 n is the number of times it's compounded since it's compounded yearly or annually it's 1 t is the time in years 5 years \[3,000(1+0.02)^{5} = $3,312~Mid~CityBank \] \[Crestline~Bank = $3,450\] seeing these two options which one has the better interest?

OpenStudy (greatlife44):

the question is asking for the difference in deposits for the two

OpenStudy (yinshy):

D